OpenStudy (anonymous):

in triangle ABC ,AB=3 and AC=9. find m angle B to the nearest degree. a. 72 b. 52 c. 18 d. 58

OpenStudy (anonymous):



OpenStudy (anonymous):

well do you know how to find BC?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

use Pythagorean theorem

OpenStudy (anonymous):

AB^2 + AC^2 = BC^2

OpenStudy (anonymous):

i tried. it want be write

OpenStudy (anonymous):

right*

OpenStudy (anonymous):

well 3^2 + 9^2 = 9 + 81 = 90 sqr 90 is about 9.48